OpenStudy (anonymous):

random question about gravity .. why don't we fall together when we come closer to each other ? I've been reading about gravitational pull and planets and it doesn't make sense how a simple change in their positions could change a lot but with us we can do what ever I get that the planets' masses are obviously ... bigger but does that make the effect of gravity on our masses that negligible. or is there some sort of force that opposes it ? sorry for the rambling

OpenStudy (mrnood):

This is simply about masses. The force of gravity between you and any other object is directly proportional to mass of that object (It is also inversely proportional to the square of the distance between you) SO Mass of earth = 6 *10^24 kg Mass of a light person = 6 * 10 kg SO the effect of the mass of the earth is 10^23 times more than the effect of the mass of your friend. Now the radius of earth is approx. 6 * 10^6m (and that's how far you are from centre of gravity!) And say your friend is 1m away So the force due to the earth is 10^-12 LESS due to its distance (squared) So the resultant is that the force due to the earth is 10^23/10^12 times more powerful i.e. 10^11 or 100 BILLION times more. =Since we know the force due to the earth is your weight (lets say 100kg!!) then the force due to your friend is 1 billionth of 1kg No wonder you don't feel the attraction! (VERY rough estimates made above!)

OpenStudy (mrnood):

The equation relating this is \[F = \frac{ GM _{1}M _{2} }{r ^{2} }\] Where F is the gravitational force between two bodies of masses M1 and M2, and r is the distance betweent hteir centres of gravity G is the Universal Gravitational Constant 6.673 84 x 10-11 m3 kg-1 s-2

OpenStudy (anonymous):

oh wow ! yeah that makes sense, thank you !